karlieeeeeee:

s (1, 10) a solution to this system of equations? y= x+5 y= 7x+3

AZ:

Hey so if it's a solution that means when we plug in x = 1 and y = 10 into BOTH of the equations, it should be true

quasia:

Hi ~ Karlie ! y = 5x 2x + y = 14 Let's take the second equation and isolate y... 2x + y = 14 Let's subtract 2x from both sides... -2x+2x+y=-2x+14 SECOND EQUATION...y=-2x+14 FIRST EQUATION.......y=5x Let's subtract the first equation from the second... 0=-7x+14 Let's add 7x to both sides... 7x+0=7x-7x+14 7x=14 x=2 Let's plug-in x=2 into the first equation.. y=5x y=(5)(2) y=10 Let's plug-in x=2 into the second equation to verify y=10... y=-2x+14 10=[(-2)(2)]+14 10=-4+14 10=10 The two lines meet at (2,10). Your question asked if these lines meet at (10,2) thus rendering such as a "solution". The answer is NO.

AZ:

So for the first equation y = x + 5 we're plugging in (1, 10) that means x = 1 and y = 10 put the numbers in 10 = 1 + 5 what is 1 + 5 =? does it equal to 10?